A manufacturer knows that their items have a normally distributed lifespan, with a mean of 13 years, and standard deviation of 4.1 years. If you randomly purchase one item, what is the probability it will last longer than 23 years?
Answer #1

Given,

= 13 , = 4.1

We convert this to standard normal as

P(X < x) = P(Z < ( x - ) / )

So,

P(X > 23) = P(Z > ( 23 - 13) / 4.1)

= P(Z > 2.44)

= 0.0073